# Some Aspects of Three-Quark Potentials (Part II)

### v2

@@ Appendix E Some details on the three-quark potential for the collinear geometry
$\eta^{1}=\frac{1+\cos^{-\frac{1}{2}}\alpha_{3}\,I\bigl{(}\cos^{2}\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr{)}}{\cos^{-\frac{1}{2}}\alpha_{2}\bigl{(}1+I\bigl{(}\sin^{2}\alpha_{2},\frac{1}{2},\tfrac{3}{4}\bigr{)}\bigr{)}-\cos^{-\frac{1}{2}}\alpha_{3}\,I\bigl{(}\cos^{2}\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr{)}}\,,$ (E.2)
@@ F.1 The limiting values of $v$
As explained in Sec.V, the limiting values of the parameter $v$ denoted as ${\text{v}}_{\text{\tiny 1}}$, $-{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$, and ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ are the solutions of Eqs.(3.12), (3.17), and (3.27), respectively. Here we examine how these solutions behave in the interval $[0,\pi]$.
@@
Figure 23: The solutions ${\text{v}}_{\text{\tiny 1}}$, $-{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$, and ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ versus $\theta_{3}$. The dotted curves correspond to $-{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$. Here $\mathsf{k}=-0.083$ (left) and $\mathsf{k}=-0.102$ (right).
The solution ${\text{v}}_{\text{\tiny 1}}$ exists and equals $0.978$. It is defined over the interval $0\leq\theta_{3}<\frac{2}{3}\pi$ and, as argued in Sec.V, is meaningful only if the triangle contains no angle greater than $\frac{2}{3}\pi$. In contrast, the solution $-{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ depends on $\theta_{3}$. It obeys $\lambda_{3}({\text{v}}_{\text{\tiny 1}}^{\text{\small-}})\leq 1$ and exists only within the narrow interval $0.66646\,\pi\leq\theta_{3}\leq\frac{2}{3}\pi$. The upper bound corresponds to $\lambda_{3}=1$, where ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}={\text{v}}_{\text{\tiny 1}}$, and the lower to $\alpha_{3}=0$, where ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}=0.967$. Recall that the condition $\alpha_{3}(v)=0$ describes a transition point between configurations II and III. The analysis of Eq.(3.27) is slightly more subtle. It shows that this Equation has two solutions within the narrow interval $0.66638\pi<\theta_{3}\leq 0.66646\pi$, and a single one for $0.66646\pi<\theta_{3}\leq\pi$. To complete the picture, we repeat the analysis for $\mathsf{k}=-0.102$  [^25] This value is motivated by phenomenology, specifically by the relation $E_{\text{\tiny QQ}}=\tfrac{1}{2}E_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ at small quark separations. with the results shown in the Figure on the right. Clearly, there are no significant differences between the two cases.
@@ 1 Example I
which follow from the sine and cosine rules. Here the notation $L(i,j|k)=|YQ_{i}|^{2}+|YQ_{j}|^{2}-2\cos\gamma_{k}|YQ_{i}||YQ_{j}|$ is used as shorthand for the denominators. We now take the IR limit keeping fixed all the $\theta$’s and letting all $\lambda_{i}\rightarrow 1$, such that all the strings in question become infinitely long. In this limit, the leading contributions to $|YQ_{i}|$ and $\gamma_{i}$ are given by $-\frac{1}{\sqrt{\mathsf{s}}}\ln(1-\lambda_{i})$ and $\frac{2}{3}\pi$, respectively, as discussed in Sec.V. As a result, (F.1) reduces to

### v3

@@ Appendix E Some details on the three-quark potential for the collinear geometry
$\eta^{1}=\frac{1+\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{3}\,I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr)}{\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{2}\bigl(1+I\bigl(\sin^{2}\hskip-2.56073pt\alpha_{2},\frac{1}{2},\tfrac{3}{4}\bigr)\bigr)-\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{3}\,I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr)}\,$, (E.2)
@@ F.1 The limiting values of $v$
As explained in Sec.V, the limiting values of the parameter $v$ denoted as ${\text{v}}_{\text{\tiny 1}}$, ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$, and ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ are the solutions of Eqs.(3.12), (3.17), and (3.27), respectively. Here we examine how these solutions behave in the interval $0,\pi]$.
@@
Figure 23: The solutions ${\text{v}}_{\text{\tiny 1}}$, ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$, and ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ versus $\theta_{3}$. The dotted curves correspond to ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$. Here $\mathsf{k}=-0.083$ (left) and $\mathsf{k}=-0.102$ (right).
The solution ${\text{v}}_{\text{\tiny 1}}$ exists and equals $0.978$. It is defined over the interval $0\leq\theta_{3}<\frac{2}{3}\pi$ and, as argued in Sec.V, is meaningful only if the triangle contains no angle greater than $\frac{2}{3}\pi$. In contrast, the solution ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ depends on $\theta_{3}$. It obeys $\lambda_{3}({\text{v}}_{\text{\tiny 1}}^{\text{\small-}})\leq 1$ and exists only within the narrow interval $0.66646\,\pi\leq\theta_{3}\leq\frac{2}{3}\pi$. The upper bound corresponds to $\lambda_{3}=1$, where ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}={\text{v}}_{\text{\tiny 1}}$, and the lower to $\alpha_{3}=0$, where ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}=0.967$. Recall that the condition $\alpha_{3}(v)=0$ describes a transition point between configurations II and III. The analysis of Eq.(3.27) is slightly more subtle. It shows that this Equation has two solutions within the narrow interval $0.66638\pi<\theta_{3}\leq 0.66646\pi$, and a single one for $0.66646\pi<\theta_{3}\leq\pi$. To complete the picture, we repeat the analysis for $\mathsf{k}=-0.102$  [^25] This value is motivated by phenomenology, specifically by the relation $E_{\text{\tiny QQ}}=\tfrac{1}{2}E_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ at small quark separations. with the results shown in the Figure on the right. Clearly, there are no significant differences between the two cases.
@@ 1 Example I
which follow from the sine and cosine rules. Here the notation $L(i,j|k)=|YQ_{i}|^{2}+|YQ_{j}|^{2}-2\cos\gamma_{k}|YQ_{i}||YQ_{j}|$ is used as shorthand for the denominators. We now take the IR limit keeping fixed all the $\theta$’s and letting all $\lambda_{i}\rightarrow 1$, such that all the strings in question become infinitely long. In this limit, the leading contributions to $|YQ_{i}|$ and $\gamma_{i}$ are given by $\frac{1}{\sqrt{\mathsf{s}}}\ln(1-\lambda_{i})$ and $\frac{2}{3}\pi$, respectively, as discussed in Sec.V. As a result, (F.1) reduces to